OpenStudy (barrelracer011):

Help will medal!!!! Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and three inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Kevin needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 3 feet 6 feet 8 feet 10 feet

OpenStudy (barrelracer011):

@texaschic101

OpenStudy (barrelracer011):

@karategirl2002

OpenStudy (karategirl2002):

here

OpenStudy (karategirl2002):

@SolomonZelman

OpenStudy (barrelracer011):

Thanks

OpenStudy (karategirl2002):

@phi

OpenStudy (karategirl2002):

@hoblos

OpenStudy (phi):

How about figuring out how many cubic yards of concrete you can buy? can you do that ?

OpenStudy (karategirl2002):

dang it i was going to say that lololololol

OpenStudy (karategirl2002):

@NooriK.

OpenStudy (barrelracer011):

I have no clue what a cubic yard is. @phi

OpenStudy (karategirl2002):

look it up

OpenStudy (karategirl2002):

be right back (BRB)

OpenStudy (barrelracer011):

I did but im still confused on how many feet a cubic yard is

OpenStudy (phi):

two things: 1) to find out how much you can buy, you don't need to know what it is. 2) it is a cube with sides all equal to 1 yard (3 ft x 3 ft x 3 ft )

OpenStudy (barrelracer011):

So it would be B 6 feet

OpenStudy (phi):

I don't know off hand. You "have to work the numbers"

OpenStudy (phi):

how many cubic yards of concrete you can buy?

OpenStudy (barrelracer011):

I cant buy a whole cubic yard because a full one is 98 dollars and I only spent 72.59

OpenStudy (phi):

yes, so you get a fraction of a whole cubic yard. what do you get ?

OpenStudy (barrelracer011):

@NooriK. So if I do 72.59/98 I will get the volume he needs?

OpenStudy (noorik.):

sorry the equation i wrote was wrong the right equation is 72.59/98= v/1 cubic yard and yes you will get the volume

OpenStudy (barrelracer011):

I did it and I got 0.7407

OpenStudy (noorik.):

yes that's the volume in cubic yards do you know the law of volume?

OpenStudy (barrelracer011):

Yes V=LxWxH

OpenStudy (phi):

next, I would change to feet (because the answer choices are in feet) multiply your fraction of a cubic yard by 27 (3*3*3 feet is 27cu ft)

OpenStudy (noorik.):

ok now you have to transform the other dimension into yards L = 10 feet = ? yards H = 3 inchs = ? feet = ? yards get it?

OpenStudy (barrelracer011):

The length would be 3.5 and 3 inches is the height

OpenStudy (barrelracer011):

3.5 yards

OpenStudy (barrelracer011):

So would it be A?

OpenStudy (noorik.):

ok, and the height is =0.085 yards

OpenStudy (barrelracer011):

.765

OpenStudy (noorik.):

then you substitute in the main law V = L * W *H

OpenStudy (noorik.):

you got the volume and the height and the length

OpenStudy (barrelracer011):

yes I would multiply 3*0.085*3

OpenStudy (noorik.):

no you would get 0.7407 = 3.5 * 0.085 * W you multiply 3.5 * 0.085 and then divide 0.7407/(3.5*0.085)

OpenStudy (barrelracer011):

I got 2.48

OpenStudy (noorik.):

that's the width in yards, you have to make it in feet

OpenStudy (barrelracer011):

10 feet

OpenStudy (noorik.):

no you would get 7.44, due to uncertainty in division and multiplying, we choose the closest answer which is 8 feet

OpenStudy (barrelracer011):

Oo ok I got it thank you so much for your help and spending your time to help me!!!

OpenStudy (noorik.):

u r welcome :)

OpenStudy (phi):

working in feet: \[ \frac{ 72.59}{98}\cdot 27 = 19.999 \ cu ft \] round to 20 cu ft of concrete The volume of the walk is 10 feet long x 3 inches high. x W ft (W is the unknown width) 10 ft x 0.25 ft x W = 20 or W= 20/2.5 = 8 ft

OpenStudy (anonymous):

It would be 8

OpenStudy (anonymous):

ft